ON Problem From Appendix C material) The exports and imports of country A in year 1 are listed below These are the only goods traded by country A Good Value of Exports Value of Imports $100 $20 300 80 100 300 5 $400 caiculate country As index ot intra-industy trade for year
Answer #1

Formula of Intra Industry Trade :

Grubel Llyod Index : 1-( Export of Industry A- Import of Industry A))/ (Export of Industry A+ Import of Industry B)

Good T =1-(100-20)/(100+20) =1-(80/120)=0.33

Good X=1-(300-80)/(300+80)=1-(220/380)=0.43

Good Y=1-(100-300)/(100+300)=1-(200/400)=0.5